OpenStudy (anonymous):

Integrate: \[\int_0^2\sqrt{x+\sqrt{x+\sqrt{x+\cdots}}}~dx\]

OpenStudy (anonymous):

Source: http://math.mit.edu/~sswatson/pdfs/qualifying_round_2014.pdf

OpenStudy (cwrw238):

sorry - i've no idea how to do this one

OpenStudy (anonymous):

No worries, it's intended to be challenging ;)

OpenStudy (kainui):

That was fun =)

ganeshie8 (ganeshie8):

\[\int_0^2\sqrt{x+\sqrt{x+\sqrt{x+\cdots}}}~dx = \int_0^2 \frac{1+\sqrt{1+4x}}{2}~dx\]

OpenStudy (anonymous):

let \[\sqrt{x+\sqrt{x+\sqrt{x...}}}=y\] \[\sqrt{x+y}=y\] \[x+y=y^2,y^2-y-x=0,y=\frac{ 1\pm \sqrt{1+4x} }{ 2 }\]

OpenStudy (anonymous):

as y is positive hence \[y=\frac{ 1+\sqrt{1+4x} }{ 2 }\]

OpenStudy (cwrw238):

tha'ts really clever

OpenStudy (anonymous):

Nice work guys!

OpenStudy (cwrw238):

a brilliant substitution!

OpenStudy (cwrw238):

That was a piece of beautiful music!!

ganeshie8 (ganeshie8):

:) cwrw you might be interested in ramanujan nested radicals, look it up in google when free

ganeshie8 (ganeshie8):

@cwrw238

OpenStudy (kainui):

Here's a related one, \[\Large \int\limits_0^1 \cos(\cos(\cos(...)))dx\]

OpenStudy (cwrw238):

@ganeshie8 thanks I will

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/DottieNumber.html

OpenStudy (anonymous):

I expected a challenging one , but this was simple

OpenStudy (anonymous):

Maybe because i have solved nested radicals like these too many times